A block of mass m_1 = 1.80 kg moving at v_1 = 2.00
Answer #1

momentum conservation suggests that

\frac{v_{2}}{v_{1}}= \frac{m_{1}}{m_{1}+m_{2}} = 0.857

\frac{v_{3}}{v_{1}}= \frac{m_{1}}{m_{1}+m_{2}+m_{3}} = 0.419